OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (anonymous):

If it was me, I would multiply the second equation by 2 and add it to the first equation ((a) because it will get rid of z by simple addition (b) it's easier when one of the equations = 0) to get 3x + 3y = 24 -> x + y = 8 Now we need another equation in x and y so we will have simultaneous equations, so in the same way, we multiply the second equation by 3 and add it to the third equation to get 5x + 11 y = 46 and you can do the rest, I think.